# Score identity Distillation: Exponentially Fast Distillation of Pretrained Diffusion Models for One-Step Generation

### v1

@@ 4.2 Loss Approximation based on Identities 1 and 2
To estimate the score $\nabla_{\bm{x}_{t}}\ln p_{\theta}(\bm{x}_{t})$, an initial thought would be to adopt a deep neural network-based approximation $f_{\psi}(\bm{x}_{t},t)\approx\mathbb{E}[\bm{x}_{g}\,|\,\bm{x}_{t}]$ by (5), which can be trained with the usual diffusion or denoising score-matching loss as
$\min_{\psi}\mathbb{E}_{q(\bm{x}_{t}\,|\,\bm{x}_{g},t)p_{\theta}(\bm{x}_{g})}[{\gamma(t)}\|f_{\psi}(\bm{x}_{t},t)-\bm{x}_{g}\|_{2}^{2}],$
where the timestep distribution $t\sim p(t)$ and weighting function $\gamma(t)$ can be defined as in Karras et al. (2022). Assuming $\bm{x}_{t}\sim q(\bm{x}_{t}\,|\,\bm{x}_{g}),\,\bm{x}_{g}=G_{\theta}(\bm{z}),\,\bm{z}\sim p(\bm{z})$, the optimal solution $\psi^{*}(\theta)$, which depends on the generator distribution determined by $\theta$, satisfies
$f_{\psi^{*}(\theta)}(\bm{x}_{t},t)=\mathbb{E}[\bm{x}_{g}\,|\,\bm{x}_{t}]=\bm{x}_{t}+\sigma_{t}^{2}\nabla_{\bm{x}_{t}}\ln p_{\theta}(\bm{x}_{t})$
@@
$\delta_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})=\sigma_{t}^{-2}(f_{\phi}(\bm{x}_{t},t)-f_{\psi^{*}(\theta)}(\bm{x}_{t},t)).$ (11)
@@
It is tempting to estimate the score difference $\delta_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})$ using an approximated score difference defined as
${\delta}_{\phi,\psi}(\bm{x}_{t}):=\sigma_{t}^{-2}(f_{\phi}(\bm{x}_{t},t)-f_{\psi}(\bm{x}_{t},t)),$ (12)
@@
${\mathcal{L}}_{\theta}\approx{\mathcal{L}}_{\theta}^{(1)}:=\mathbb{E}_{\bm{x}_{t}\sim p_{\theta}(\bm{x}_{t})}\left[\|{\delta}_{\phi,\psi}(\bm{x}_{t})\|_{2}^{2}\right].$ (13)
@@
$\triangle_{\psi,\psi^{*}(\theta)}(\bm{x}_{t}):=\sigma_{t}^{-2}({f_{\psi}(\bm{x}_{t},t)-f_{\psi^{*}(\theta)}(\bm{x}_{t},t)}),$ (14)
we have ${\delta}_{\phi,\psi}(\bm{x}_{t})={\delta}_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})-\triangle_{\psi,\psi^{*}(\theta)}(\bm{x}_{t})$ and
${\mathcal{L}}_{\theta}^{(1)}={\mathcal{L}}_{\theta}\,+\,$ $\mathbb{E}_{p_{\theta}(\bm{x}_{t})}\big[\|\triangle_{\psi,\psi^{*}(\theta)}(\bm{x}_{t})\|_{2}^{2}$
$-2\triangle_{\psi,\psi^{*}(\theta)}(\bm{x}_{t})^{T}\delta_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})\big].$ (15)
Therefore, how well ${\mathcal{L}}_{\theta}^{(1)}$ approximates the true loss ${\mathcal{L}}_{\theta}$ heavily depends on not only the score approximation error $\triangle_{\psi,\psi^{*}(\theta)}(\bm{x}_{t})$ but also the score difference $\delta_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})$. For a given $\theta$, although one can control $\triangle_{\psi,\psi^{*}(\theta)}(\bm{x}_{t})$ by minimizing (10), it would be difficult to ensure that influence of the score difference $\delta_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})$ would not dominate the true loss ${\mathcal{L}}_{\theta}$, especially during the intial phase of training when $p_{\theta}(\bm{x}_{t})$ does not match $p_{\text{data}}(\bm{x}_{t})$ well.
@@
$\hat{{\mathcal{L}}}_{\theta}^{(1)}=\|{\delta}_{\phi,\psi}(\bm{x}_{t})\|_{2}^{2},\penalty 10000\$(16)
$\bm{x}_{t}=\bm{x}_{g}+\sigma_{t}\bm{\epsilon}_{t},\,\,\bm{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (17)
$\bm{x}_{g}=G_{\theta}(\sigma_{\text{init}}\bm{z}),\penalty 10000\ \bm{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (18)
@@
The proof is presented in Appendix˜B. The example in this proposition shows that although minimizing the objective $\mathcal{L}_{\theta}$ leads to the optimal generator parameter $\theta^{*}=0$, the loss $\hat{\mathcal{L}}_{\theta}^{(1)}$ would provide no meaningful gradient towards $\theta^{*}$.
@@ 4.3 Loss Approximation via Projected Score Matching
${\mathcal{L}}_{\theta}=\mathbb{E}_{q(\bm{x}_{t}\,|\,\bm{x}_{g},t)p_{\theta}(\bm{x}_{g})}\left[\sigma_{t}^{-2}{\delta_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})^{T}(f_{\phi}(\bm{x}_{t},t)-\bm{x}_{g})}\right]$.
The proof, based on Identity 3, is deferred to Appendix˜A. We approximate the loss by substituting $\psi^{*}(\theta)$ in (19) with its approximation $\psi$, leading to an approximated loss ${\mathcal{L}}_{\theta}^{(2)}$ as
@@
$\mathbb{E}_{q(\bm{x}_{t}\,|\,\bm{x}_{g},t)p_{\theta}(\bm{x}_{g})}\left[\sigma_{t}^{-2}{\delta_{\phi,\psi}(\bm{x}_{t})^{T}(f_{\phi}(\bm{x}_{t},t)-\bm{x}_{g})}\right]$
$\!\!={\mathcal{L}}_{\theta}-\,\mathbb{E}_{q(\bm{x}_{t}\,|\,\bm{x}_{g},t)p_{\theta}(\bm{x}_{g})}$
$\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \left[\sigma_{t}^{-2}{\triangle_{\psi,\psi^{*}(\theta)}(\bm{x}_{t})^{T}(f_{\phi}(\bm{x}_{t},t)-\bm{x}_{g})}\right].$ (20)
Comparing (20) to (15) indicates that ${\mathcal{L}}_{\theta}^{(2)}$ is directly influenced by neither the norm $\|\triangle_{\psi,\psi^{*}(\theta)}(\bm{x}_{t})\|_{2}^{2}$ nor the score difference $\delta_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})$ given by (11). Initially in training, the discrepancy between the estimated and actual scores for the generator distribution may amplify the value of $\triangle_{\psi,\psi^{*}(\theta)}(\bm{x}_{t})$, whereas the difference between the pre-trained score for the real data distribution and the actual score for the generator distribution may inflate $\delta_{\phi,\psi^{*}(\theta)}(\bm{x}_{t})$. By contrast, the term $f_{\phi}(\bm{x}_{t},t)-\bm{x}_{g}$ within (20) reflects the efficacy of the pre-trained model in denoising corrupted fake data, which tends to be more stable.
@@
Under the setting of Proposition˜4, the gradient of loss ${\mathcal{L}}_{\theta}^{(2)}$ can be estimated as
$\nabla_{\theta}\hat{L}_{\theta}^{(2)}=-({1+\sigma_{t}^{2}})^{-1}{\delta}_{\phi,\psi}(\bm{x}_{t})\nabla_{\theta}G_{\theta}(\sigma_{\text{init}}\bm{z}),$
@@
which involves the product of the approximated score difference ${\delta}_{\phi,\psi}(\bm{x}_{t})=-\frac{\psi}{1+\sigma_{t}^{2}}$ and the gradient of the generator.
@@
We note the product of the approximated score difference and $\nabla_{\theta}G_{\theta}(\sigma_{\text{init}}\bm{z})$ is used to construct the loss for Diff-Instruct (Luo et al., 2023c), which has been shown to be able to distill a pretrained diffusion model with satisfactory performance. Thus for the toy example where ${\mathcal{L}}_{\theta}^{(1)}$ fails, using ${\mathcal{L}}_{\theta}^{(2)}$ can provide useful gradient to guide the generator.

### v2

@@ 4.2 Loss Approximation based on Identities 1 and 2
To estimate the score $\nabla_{\boldsymbol{x}_{t}}\ln p_{\theta}(\boldsymbol{x}_{t})$, an initial thought would be to adopt a deep neural network-based approximation $f_{\psi}(\boldsymbol{x}_{t},t)\approx\mathbb{E}[\boldsymbol{x}_{g}\,|\,\boldsymbol{x}_{t}]$ by (5), which can be trained with the usual diffusion or denoising score-matching loss as
$\min_{\psi}\mathbb{E}_{q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g},t)p_{\theta}(\boldsymbol{x}_{g})}[{\gamma(t)}\|f_{\psi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g}\|_{2}^{2}],$
where the timestep distribution $t\sim p(t)$ and weighting function $\gamma(t)$ can be defined as in Karras et al. (2022). Assuming $\boldsymbol{x}_{t}\sim q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g}),\,\boldsymbol{x}_{g}=G_{\theta}(\boldsymbol{z}),\,\boldsymbol{z}\sim p(\boldsymbol{z})$, the optimal solution $\psi^{*}(\theta)$, which depends on the generator distribution determined by $\theta$, satisfies
$f_{\psi^{*}(\theta)}(\boldsymbol{x}_{t},t)=\mathbb{E}[\boldsymbol{x}_{g}\,|\,\boldsymbol{x}_{t}]=\boldsymbol{x}_{t}+\sigma_{t}^{2}\nabla_{\boldsymbol{x}_{t}}\ln p_{\theta}(\boldsymbol{x}_{t})$
@@
$\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})=\sigma_{t}^{-2}(f_{\phi}(\boldsymbol{x}_{t},t)-f_{\psi^{*}(\theta)}(\boldsymbol{x}_{t},t)).$ (11)
@@
It is tempting to estimate the score difference $\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})$ using an approximated score difference defined as
${\delta}_{\phi,\psi}(\boldsymbol{x}_{t}):=\sigma_{t}^{-2}(f_{\phi}(\boldsymbol{x}_{t},t)-f_{\psi}(\boldsymbol{x}_{t},t)),$ (12)
@@
${\mathcal{L}}_{\theta}\approx{\mathcal{L}}_{\theta}^{(1)}:=\mathbb{E}_{\boldsymbol{x}_{t}\sim p_{\theta}(\boldsymbol{x}_{t})}\left[\|{\delta}_{\phi,\psi}(\boldsymbol{x}_{t})\|_{2}^{2}\right].$ (13)
@@
$\triangle_{\psi,\psi^{*}(\theta)}(\boldsymbol{x}_{t}):=\sigma_{t}^{-2}({f_{\psi}(\boldsymbol{x}_{t},t)-f_{\psi^{*}(\theta)}(\boldsymbol{x}_{t},t)}),$ (14)
we have ${\delta}_{\phi,\psi}(\boldsymbol{x}_{t})={\delta}_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})-\triangle_{\psi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})$ and
${\mathcal{L}}_{\theta}^{(1)}={\mathcal{L}}_{\theta}\,+\,$ $\mathbb{E}_{p_{\theta}(\boldsymbol{x}_{t})}\big[\|\triangle_{\psi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})\|_{2}^{2}$
$-2\triangle_{\psi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})^{T}\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})\big].$ (15)
Therefore, how well ${\mathcal{L}}_{\theta}^{(1)}$ approximates the true loss ${\mathcal{L}}_{\theta}$ heavily depends on not only the score approximation error $\triangle_{\psi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})$ but also the score difference $\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})$. For a given $\theta$, although one can control $\triangle_{\psi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})$ by minimizing (10), it would be difficult to ensure that influence of the score difference $\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})$ would not dominate the true loss ${\mathcal{L}}_{\theta}$, especially during the intial phase of training when $p_{\theta}(\boldsymbol{x}_{t})$ does not match $p_{\text{data}}(\boldsymbol{x}_{t})$ well.
@@
$\hat{{\mathcal{L}}}_{\theta}^{(1)}=\|{\delta}_{\phi,\psi}(\boldsymbol{x}_{t})\|_{2}^{2},$~ (16)
$\boldsymbol{x}_{t}=\boldsymbol{x}_{g}+\sigma_{t}\boldsymbol{\epsilon}_{t},\,\,\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (17)
$\boldsymbol{x}_{g}=G_{\theta}(\sigma_{\text{init}}\boldsymbol{z}),~\boldsymbol{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (18)
@@
The proof is presented in Appendix D. The example in this proposition shows that although minimizing the objective $\mathcal{L}_{\theta}$ leads to the optimal generator parameter $\theta^{*}=0$, the loss $\hat{\mathcal{L}}_{\theta}^{(1)}$ would provide no meaningful gradient towards $\theta^{*}$.
@@ 4.3 Loss Approximation via Projected Score Matching
${\mathcal{L}}_{\theta}=\mathbb{E}_{q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g},t)p_{\theta}(\boldsymbol{x}_{g})}\left[\sigma_{t}^{-2}{\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})^{T}(f_{\phi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g})}\right]$.
The proof, based on Identity 3, is deferred to Appendix C. We approximate the loss by substituting $\psi^{*}(\theta)$ in (19) with its approximation $\psi$, leading to an approximated loss ${\mathcal{L}}_{\theta}^{(2)}$ as
@@
$\mathbb{E}_{q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g},t)p_{\theta}(\boldsymbol{x}_{g})}\left[\sigma_{t}^{-2}{\delta_{\phi,\psi}(\boldsymbol{x}_{t})^{T}(f_{\phi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g})}\right]$
$\!\!={\mathcal{L}}_{\theta}-\,\mathbb{E}_{q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g},t)p_{\theta}(\boldsymbol{x}_{g})}$
$~~~~~~~~~~\left[\sigma_{t}^{-2}{\triangle_{\psi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})^{T}(f_{\phi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g})}\right].$ (20)
Comparing (20) to (15) indicates that ${\mathcal{L}}_{\theta}^{(2)}$ is directly influenced by neither the norm $\|\triangle_{\psi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})\|_{2}^{2}$ nor the score difference $\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})$ given by (11). Initially in training, the discrepancy between the estimated and actual scores for the generator distribution may amplify the value of $\triangle_{\psi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})$, whereas the difference between the pre-trained score for the real data distribution and the actual score for the generator distribution may inflate $\delta_{\phi,\psi^{*}(\theta)}(\boldsymbol{x}_{t})$. By contrast, the term $f_{\phi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g}$ within (20) reflects the efficacy of the pre-trained model in denoising corrupted fake data, which tends to be more stable.
@@
Under the setting of Proposition 4, the gradient of loss ${\mathcal{L}}_{\theta}^{(2)}$ can be estimated as
$\nabla_{\theta}\hat{L}_{\theta}^{(2)}=-({1+\sigma_{t}^{2}})^{-1}{\delta}_{\phi,\psi}(\boldsymbol{x}_{t})\nabla_{\theta}G_{\theta}(\sigma_{\text{init}}\boldsymbol{z}),$
@@
which involves the product of the approximated score difference ${\delta}_{\phi,\psi}(\boldsymbol{x}_{t})=-\frac{\psi}{1+\sigma_{t}^{2}}$ and the gradient of the generator.
@@
We note the product of the approximated score difference and $\nabla_{\theta}G_{\theta}(\sigma_{\text{init}}\boldsymbol{z})$ is used to construct the loss for Diff-Instruct (Luo et al., 2023c), which has been shown to be able to distill a pretrained diffusion model with satisfactory performance. Thus for the toy example where ${\mathcal{L}}_{\theta}^{(1)}$ fails, using ${\mathcal{L}}_{\theta}^{(2)}$ can provide useful gradient to guide the generator.
4.4 Fused Loss of SiD
${\mathcal{L}}_{\theta}^{(2)}=$~ ${\mathcal{L}}_{\theta}^{(1)}+\mathbb{E}_{\boldsymbol{x}_{g}\sim p_{\theta}(\boldsymbol{x}_{g})}\mathbb{E}_{\boldsymbol{x}_{t}\sim q(\boldsymbol{x}_{t}\,|\,\boldsymbol{x}_{g},t)}$
$\left[{\sigma_{t}^{-2}}{{\delta}_{\phi,\psi}(\boldsymbol{x}_{t})^{T}(f_{\psi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g})}\right].$ (21)
$\hat{{\mathcal{L}}}_{\theta}^{\triangle}=\hat{{\mathcal{L}}}_{\theta}^{(2)}-\hat{{\mathcal{L}}}_{\theta}^{(1)}={\sigma_{t}^{-2}}{\delta}_{\phi,\psi}(\boldsymbol{x}_{t})^{T}(f_{\psi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g})$
directs the gradient towards the desired direction.
Thus we are propelled to consider the loss
In summary, the weighted loss is expressed as
$\textstyle\tilde{L}_{\theta}(\boldsymbol{x}_{t},t,\phi,\psi)=(1-\alpha)\frac{\omega(t)}{{\sigma_{t}^{4}}}\|f_{\phi}(\boldsymbol{x}_{t},t)-f_{\psi}(\boldsymbol{x}_{t},t)\|_{2}^{2}$
$\textstyle+\frac{\omega(t)}{{\sigma_{t}^{4}}}(f_{\phi}(\boldsymbol{x}_{t},t)-f_{\psi}(\boldsymbol{x}_{t},t))^{T}(f_{\psi}(\boldsymbol{x}_{t},t)-\boldsymbol{x}_{g}),$ (23)
where $\boldsymbol{x}_{t}$ is generated as in (18) and $\omega(t)$ are weighted coefficients that need to be specified. To compute the gradient of the above equation, SiD backpropagates the gradient through both $\phi$ and $\psi$ by calculating two score gradients ($i.e.$, gradients of scores) as
$\textstyle\nabla_{\theta}f_{\phi}(\boldsymbol{x}_{t},t)=\frac{\partial f_{\phi}(\boldsymbol{x}_{t},t)}{\partial\boldsymbol{x}_{t}}\nabla_{\theta}G_{\theta}(\sigma_{\text{init}}\boldsymbol{z})$ (24)
$\textstyle\nabla_{\theta}f_{\psi}(\boldsymbol{x}_{t},t)=\frac{\partial f_{\psi}(\boldsymbol{x}_{t},t)}{\partial\boldsymbol{x}_{t}}\nabla_{\theta}G_{\theta}(\sigma_{\text{init}}\boldsymbol{z}).$
This feature distinguishes SiD from Diff-Instruct and DMD that do not use score gradients $\frac{\partial f_{\phi}(\mathbf{x}_{t},t)}{\partial\boldsymbol{x}_{t}}$ and $\frac{\partial f_{\psi}(\boldsymbol{x}_{t},t)}{\partial\boldsymbol{x}_{t}}$.
4.5 Noise Weighting and Scheduling
$\omega(t)=C\sigma_{t}^{4}/\|\boldsymbol{x}_{g}-f_{\phi}(\boldsymbol{x}_{t},t)\|_{1,sg}~~.$ (25)
The distillation process is outlined in Algorithm 1. The one-step generation procedure is straightforward: $\boldsymbol{x}=G_{\theta}(\sigma_{\text{init}}\boldsymbol{z}),~\boldsymbol{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ where $\sigma_{\text{init}}$, by default set to 2.5, remains consistent throughout distillation and generation.
[…]
tableComparison of various deep generative models trained on CIFAR-10 without label conditioning. The best and second-best one/few-step generators under the FID or IS metric are highlighted with bold and italic bold, respectively.
[…]
figureEvolution of FIDs for the SiD generator during the distillation of the EDM teacher model pretrained on CIFAR-10 (unconditional), using $\alpha=1.0$ or $\alpha=1.2$ and a batch size of 256. The performance of EDM, along with DMD and Diff-Instruct, is depicted with horizontal lines in purple, green, and red, respectively.
5 Experimental Results
[…]
tableAnalogous to Table 4.5 for CIFAR-10 (conditional).
[…]
figureAnalogous to Fig. 4.5 for CIFAR-10 (conditional).
Implementation details. We implement SiD based on the EDM (Karras et al., 2022) code base and we initialize both the generator $G_{\theta}$ and its score estimation network $f_{\psi}$ by copying the architecture and parameters of the pretrained score network $f_{\phi}$ from EDM (Karras et al., 2022). We provide the other implementation details in Appendix E.
[…]
tableAnalogous to Table 4.5 for ImageNet 64x64 with label conditioning. The Precision and Recall metrics are also included.
[…]
figureAnalogous plot to Fig. 4.5 for ImageNet 64x64. The batch size is 8192. See the results of batch size 1024 in Fig. E.
[…]
tableAnalogous to Table 4.5 for FFHQ 64x64.
[…]
figureAnalogous plot to Fig. 4.5 for FFHQ 64x64. The batch size is 512.
[…]
tableAnalogous to Table 4.5 for AFHQ-v2 64x64.
[…]
figureAnanagous plot to Fig. 4.5 for AFHQ-v2 64x64. The batch size is 512.
Ablation Study and Parameter Settings. We provide ablation studies and discuss parameter settings in Appendix A.
5.1 Benchmark Performance
Our comprehensive evaluation compares SiD against leading deep generative models, encompassing both distilled diffusion models and those built from scratch. Random images generated by SiD in a single step are displayed in Figs. 7-10 in the Appendix.
The comparative analysis, detailed in Tables 4.5-5 and illustrated in Figs. 4.5-5, underlines the single-step SiD generator’s proficiency in leveraging the insights from the pretrained EDM (teacher diffusion model) across a variety of benchmarks, including CIFAR-10 (both conditional and unconditional formats), ImageNet 64x64, FFHQ 64x64, and AFHQ-v2 64x64. Remarkably, the SiD-trained generator surpasses the EDM teacher in nearly all tested environments, showcasing its enhanced performance not just relative to the original multi-step teacher model but also against a broad spectrum of cutting-edge models, from traditional multi-step diffusion models to the latest single-step distilled models and GANs. The sole deviation in this pattern occurs with ImageNet 64x64, where SiD, at $\alpha=1.2$, attains an FID of 1.524, which is exceeded by Jabri et al. (2022)’s RIN at 1.23 FID with 1000 steps and the teacher model VP-EDM’s 1.36 FID with 511 steps.
Convergence Speed. Our SiD generator, designed for distilling pretrained diffusion models, rapidly achieves the capability to generate photo-realistic images in a single step. This efficiency is showcased in Fig. 1 for the EDM model pretrained on ImageNet 64x64 and in Fig. 2 for CIFAR 32x32 (unconditional). The performance of the SiD method is further highlighted in Figs. 4.5-5, where the x-axis represents the thousands of images processed during training. These figures track the FID’s evolution across four datasets for both $\alpha=1$ and $\alpha=1.2$, demonstrating a roughly linear relationship between the logarithm of the FID and the logarithm of the number of processed images. This relationship indicates that FID decreases exponentially as distillation progresses, a trend that is observed or expected to eventually slow down and approach a steady state.
6 Conclusion
Impact Statement
Acknowledgments
The authors would like to thank Dr. Zhuoran Yang and Weijian Luo for their valuable comments and suggestions.
Appendix for Score identity Distillation
Appendix A Ablation Study and Parameter Settings
The results indicate a stable performance of the model when $\alpha$ varies from 0 to 1.2. A negative value of $\alpha$ results in large FIDs. This observation supports our analysis in Section 4.2 that directly optimizing $\mathcal{L}_{\theta}^{(1)}$ given by (13) may not lead to meaningful improvement, as our loss, shown in (22), is $\mathcal{L}_{\theta}^{(2)}-\alpha\mathcal{L}_{\theta}^{(1)}$. As $\alpha$ increases within the tested range, we observe a gradual improvement in IS and FID performance, peaking at $\alpha=1$ or $\alpha=1.2$. Based on these findings, we select $\alpha=1$ or $\alpha=1.2$ for all our experiments, although a more refined grid search on $\alpha$ might reveal even better performance outcomes.
Instead, we increased the overall batch size from 1024 to 8192 (while keeping the batch per GPU unchanged at 16, requiring more gradient accumulation rounds) and reduced the learning rate from 5e-6 to 4e-6. Under $\alpha=1$, we observed stable performance, while under $\alpha=1.2$, we observed occasional spikes in FID. Upon examining the generations corresponding to these spikes, as shown in the fifth image of Fig. 5 and Fig. 6 in the Appendix, we found interesting patterns where certain uncommon features, such as nests containing birds, were exaggerated. However, with a batch size as large as 8192, these occasional spikes did not seem to significantly impact the overall declining trend, which was roughly log-log linear initially and gradually leveled off. With that said, when the batch size was reduced to 1024, the sudden divergence could potentially be caused by such a spike, as observed in Fig. E.
The drawback of using a larger batch size in this case is that it takes SiD longer to outperform Diff-instruct and DMD, as clearly shown by comparing the FID trajectories in Figs. 5 and E. Although it’s feasible to develop more advanced strategies, including progressively increasing the batch size, annealing the learning rate, and implementing gradient clipping, we’ll reserve these for future study.